OpenStudy (laddiusmaximus):

calculate the derivative tln(t^(2)+1)

OpenStudy (anonymous):

t*ln(t^(2)+1) use the product rule and then apply derivative of ln(x) and u gotta take derivative of of (t^2+1 then u get =t * 1/(t^2+1)* 2t+ln(t^(2)+1) *1

OpenStudy (laddiusmaximus):

no chain rule?

OpenStudy (anonymous):

take derivative of of (t^2+1) is the chain rule ln(x) what ever is in side the parameter u gotta take derivative too let say ln(t^(2)+1) u take derivative of ln(t^(2)+1) u get 1/t^(2)+1) times the derivative of whatever is inside parenthesisin this case t^(2)+1) = 2t